### v1

# Enhancing Peer Review in Astronomy: A Machine Learning and Optimization Approach to Reviewer Assignments for ALMA

Abstract
The increasing volume of papers and proposals undergoing peer review emphasizes the pressing need for greater automation to effectively manage the growing scale. In this study, we present the deployment and evaluation of machine learning and optimization techniques for assigning proposals to reviewers that was developed for the Atacama Large Millimeter/submillimeter Array (ALMA) during the Cycle 10 Call for Proposals issued in 2023. By utilizing topic modeling algorithms, we identify the proposal topics and assess reviewers’ expertise based on their historical ALMA proposal submissions. We then apply an adapted version of the assignment optimization algorithm from PeerReview4All (Stelmakh et al., 2021a) to maximize the alignment between proposal topics and reviewer expertise. Our evaluation shows a significant improvement in matching reviewer expertise: the median similarity score between the proposal topic and reviewer expertise increased by 51 percentage points compared to the previous cycle, and the percentage of reviewers reporting expertise in their assigned proposals rose by 20 percentage points. Furthermore, the assignment process proved highly effective in that no proposals required reassignment due to significant mismatches, resulting in a savings of 3 to 5 days of manual effort.
software: gensim (Řehůřek & Sojka, 2010), gurobi (Gurobi Optimization, LLC, 2023), matplotlib (Hunter, 2007), PeerReview4All (Stelmakh et al., 2021a), spaCy (Honnibal et al., 2020)
1 Introduction
Peer review is the cornerstone of modern scientific research, playing a crucial role in the allocation of billions in grant funding and providing access to high-cost facilities such as supercomputers and telescopes, both ground- and space-based. Traditionally, this process relies on an invited panel of experts to evaluate and recommend which proposals should be accepted. This model has been widely adopted across the scientific community, including by major astronomical observatories. One prominent example is the Atacama Large Millimeter/submillimeter Array (ALMA), which utilized panel-based peer review when it began science operations in 2011. As the largest ground-based observatory and the most sensitive telescope ever constructed for high-resolution imaging of the submillimeter sky, ALMA is in high demand, making an efficient peer review process essential to its operations.
The increasing volume of proposals submitted to major observatories has put the traditional panel-based peer review process under strain. The Hubble Space Telescope (HST), the James Web Space Telescope (JWST), the European Southern Observatory (ESO), and ALMA for example receive between 1000 to 2000 proposals per year. Managing this large volume requires an ever-growing number of reviewers, complicating the logistics of organizing panel meetings, which can require 200 or more participants. These challenges have prompted some observatories to evolve the peer review process from the traditional panel approach to allow for proposal review without a face-to-face or virtual meeting of the reviewers. HST and JWST now use external reviewers in this manner for the smaller time requests, and distributed peer review (Merrifield & Saari, 2009) is in use at Gemini for Fast Turnaround proposals (Andersen et al., 2019), the European Southern Observatory (Patat et al., 2019; Jerabkova et al., 2023), and ALMA (Donovan Meyer et al., 2022), where each proposal team designates one person to participate in the review process.
In addition to logistical challenges, the assignment of proposals to reviewers has also become increasingly complex due to the large volume of submissions. In smaller systems, this task can often be handled manually, but for the largest observatories, automated approaches are necessary. Commonly, proposals and reviewers are tagged with topical keywords to facilitate matching based on shared expertise. However, this method has limitations, particularly in the accurate assignment of keywords to both proposals and reviewers (Strolger et al., 2023). With the increasing use of distributed peer review processes, emphasis is also placed on matching the proposals to the reviewers expertise since the reviewers will not have a group discussion of the proposals. As a result, HST, JWST, and ESO have begun incorporating machine learning and optimization techniques to improve the matching of proposals to reviewer’s expertise (Strolger et al., 2017, 2023; Kerzendorf et al., 2020).
These advanced algorithms typically work in two phases (see Section 3 in Shah, 2022). First, machine learning methods are used to quantify the similarity between the subject matter of a proposal and the expertise of the reviewer. Based on these similarities, proposals are then assigned to reviewers through optimization of a pre-determined metric. Notably, automated machine learning algorithms have been integrated into the assignment processes of HST and JWST (Strolger et al., 2017, 2023) and in a pilot initiative at ESO (Patat et al., 2019; Kerzendorf et al., 2020). Following these examples, ALMA introduced machine learning and optimization techniques into its review process in 2023 (ALMA Cycle 10).
This paper details ALMA’s deployment of this new assignment process and evaluates its performance. Our main contributions are as follows:
  • Deployment. We deploy machine learning and optimization techniques to assign reviewers to the submitted proposals.
    • We use topic modeling to compute a predicted expertise (termed “similarity,” a number between 0 and 1) of each reviewer for each submitted proposal.
    • We then employ the PeerReview4All optimization algorithm (Stelmakh et al., 2021a) to optimally assign reviewers to the proposals, taking the aforementioned predicted expertise as input.
  • Evaluation. We evaluate the quality of the deployed assignment in several ways.
    • In the current cycle, assigned reviewer-paper pairs have a median similarity score of 0.71, up from 0.20 in the previous cycle.
    • We also conducted a reviewer survey, where we find that reviewers indicated a higher level of expertise in their proposal assignments with the new algorithm. The percentage of reviewers claiming expertise in their assigned proposals rose from 45% in the previous cycle to 65% in this cycle, while those indicating little or no knowledge decreased from 10% to 5%.
    • The process was highly effective, as no manual adjustments were required to address any significant mismatches between reviewer expertise and proposal topics, as compared to 3 to 5 days of manual effort previously.
The paper is organized as follows. Section 2 provides an overview of the review process used by ALMA in previous cycles and the motivation to implement a new approach. In Section 3, we describe related work in astronomy and other fields that have implemented machine learning and optimization techniques that was incorporated into ALMA’s efforts. Section 4 describes the deployment of the new system, including the use of machine learning to infer the topic of individual proposals and the reviewer’s expertise, and the assignment of proposals to individual reviewers. Section 5 evaluates the results of the new process, including back testing the results to previous cycles and presenting the results from actual use in Cycle 10. Section 6 discusses the results and summarizes the conclusions.
2 Overview of ALMA Proposal Assignment in Previous Cycles
ALMA began offering observing time in 2011 and solicits proposals annually. The number of submitted proposals increased steadily from 919 proposals in 2011 (Cycle 0) to a peak of 1836 in 2018 (Cycle 6), and currently about 1700 proposals are submitted annually. PIs submit their proposals under one of five scientific categories. [^1] Category 1: Cosmology and the high redshift universe. Category 2: Galaxies and galactic nuclei. Category 3: Interstellar medium, star formation and astrochemistry. Category 4: Circumstellar disks, exoplanets and the solar system. Category 5: Stellar evolution and the Sun. Within each category, they also select one or two keywords from a list provided by ALMA that best describe the topic of their proposal.
Initially, ALMA employed a panel-based review process, where panels of 6 to 9 experts for each of the five scientific categories evaluated around 100 proposals each. Multiple panels were formed per category as needed depending on the number of submitted proposals. However, this system became strained as the number of proposals increased, prompting ALMA to implement a distributed peer review process. This system was first used in 2019 (Cycle 7) for a “supplemental” call to fill time on the Morita Array (also known as the Atacama Compact Array) that was not allocated in the main call (Carpenter et al., 2020). ALMA used distributed peer review in the main call for the first time in 2021 (Cycle 8) for a subset of proposals, and adopted the distributed process on all proposals except for ALMA Large Programs in 2022 (Cycle 9).
In the distributed peer review system, the Principal Investigator (PI) of each proposal nominates a team member to serve as a reviewer. Each reviewer assesses 10 proposals, and each proposal receives 10 reviews. Reviewers rank the proposals from 1 (strongest) to 10 (weakest), and the final overall rankings are based on the averaged individual rankings. The process is detailed by Donovan Meyer et al. (2022). The number of proposals evaluated by this process was 1497, 1729, and 1635 in Cycles 8, 9, and 10, respectively.
Initially, ALMA used a “category” and “keyword” approach to match proposals with reviewers in the distributed system. Reviewers specified their scientific expertise using the same list of keywords used by PIs to characterize their proposals. A series of “rules” were specified that established the priority order on how to assign proposals to reviewers (see Donovan Meyer et al., 2022). The top priority was to assign proposals where one or more of the proposal keywords specified by the PI overlapped with the expertise keywords specified by the reviewer. Proposals that had the fewest available expert reviewers were assigned first. The assignment process successfully aligned the keywords of the assigned proposals to the reviewer’s expertise keywords in about 95% of cases. To gauge the success of the assignment process from the reviewer’s perspective, each reviewer had the option to rate their expertise on the proposal, with possible options of “My field of expertise”, “Some knowledge”, or “Little or no knowledge”. Reviewers indicated that had at least some knowledge in 90% of the assigned proposals and little or no knowledge in just 10% of the assignments (Donovan Meyer et al., 2022).
Despite the success of the category and keyword approach, some limitations persist. In the 10% of cases where reviewers reported little or no expertise, the proposal keywords often matched the reviewer’s expertise keywords (Donovan Meyer et al., 2022). This discrepancy could arise from several factors: the keywords may be too broad, reviewers might have selected incorrect keywords, they may be conservative in assessing their expertise, or they might be relatively inexperienced, as ALMA allows student reviewers. Additionally, keywords selected by reviewers might not fully capture their expertise, making it difficult to identify all suitable reviewers for a given proposal. Another challenge is the logistical need to assign 10 proposals to each reviewer and ensure each proposal is reviewed 10 times. The assignment algorithm had some deficiencies in that not all reviewers received proposals aligned with their expertise, and in extreme cases, some reviewers were assigned proposals outside their field (e.g., a circumstellar disk proposal assigned to an extragalactic expert). These mismatches were resolved by identifying reviewers to swap individual assignments. However, this process took 3 to 5 days of manual effort to complete, and was the most time consuming task in the assignment process.
ALMA set out to implement a new approach to assign proposals to reviewers with three objectives. First, the proposal assignment process should result in an overall increase in the expertise as measured by the similarity between the proposal topic and the reviewer’s expertise. Second, reviewers should indicate a higher level of expertise in their assignments. And finally, the process needs to be automated and robust to reduce the amount of time spent on manual re-assignments.
3 Related work
Other observatories and scientific fields face similar challenges as ALMA in that they receive hundreds if not thousands of proposals or papers that needed to be reviewed. Automated techniques are needed to assign proposals or papers to reviewers in an optimal manner. This section describes some of these previous efforts. While not an exhaustive review of the subject, we highlight previous efforts that most impacted ALMA’s implementation of machine learning and assignment optimization. We begin by discussing two prior deployments for proposal reviewing in astronomy, followed by deployments for paper review in computer science, and conclude the section by noting differences with the present work.
3.1 Space Telescope Science Institute
The Space Telescope Science Institute has developed a tool called the Proposal Auto-Categorizer and Manager for Time Allocation Review (PACMan; Strolger et al., 2017, 2023) that provides a number of services for proposal evaluation for the Hubble Space Telescope (HST) and the James Web Space Telescope (JWST). PACMan functionality includes using machine learning techniques to classify proposals into broad topics, inferring the expertise of reviewers based on their abstracts from their papers in the Astrophysics Data System (ADS), [^2] https://ui.adsabs.harvard.edu matching proposals to reviewers, and identifying conflicts of interest, among other tasks. Proposal classification and reviewer’s expertise are determined using the Term Frequency-Inverse Document Frequency (TF-IDF) model. Their results indicate that the top topical category inferred by PACMan agrees with the manual classification in 84% of the cases on average across categories. Similarly, the category of scientific expertise inferred for the reviewers has a high correspondence with the expertise provided by the reviewers.
3.2 European Southern Observatory
The European Southern Observatory (ESO) implemented machine learning techniques (Kerzendorf, 2019; Kerzendorf et al., 2020) for their pilot run of distributed peer review (Patat et al., 2019, see also Jerabkova et al. 2023), which was run in parallel with a traditional panel review. The topic of each proposal was inferred from the proposal text using the TF-IDF model, while the reviewer expertise was established based on their publication history in ADS. The dot product of the proposal vector and the reviewer expertise (“cosine distance”) then provided a measurement of the expertise. The proposals were then assigned to reviewers by optimizing the similarities of the assigned proposals. The results show a high correlation between the predicted expertise and the self-reported expertise from the reviewers (Kerzendorf et al., 2020).
3.3 Computer Science
Computer science has used distributed peer review for well over a decade to select papers submitted to conferences (for an overview of the review process, see Shah, 2022). The largest conferences can have over 13,000 submissions (NeurIPS, 2023), and therefore automated tools are essential to handle so many papers. These conferences review full papers (and not just abstracts), are frequently a terminal venue for publication, and are rated at par or better than journals.
Here is a short summary of the review process in such conferences. After papers are submitted to the conference, they are assigned to reviewers in a (semi-)automated fashion in a manner similar to that adopted at ALMA. First, a similarity score is computed via one or more of the following methods: automated text-matching (Charlin et al., 2012; Mimno & McCallum, 2007; Cohan et al., 2020), computing the intersection between reviewer- and author-chosen subject areas, and by allowing reviewers to manually specify their interest and/or expertise for various papers (called “bidding”; see Fiez et al., 2020). If more than one of these methods are used, then the outcomes are combined to obtain a single similarity score for each reviewer-paper pair. Subsequently, an optimization procedure assigns reviewers to papers. The metric that is optimized depends on the conference, and includes maximizing the total sum of similarity scores of the assignments (Goldsmith & Sloan, 2007; Charlin et al., 2012), or maximizing the similarities for the papers that are worst-off in terms of available expertise (Stelmakh et al., 2021a; Kobren et al., 2019). We refer the reader to Stelmakh et al. (2023), Saveski et al. (2023), and Stelmakh et al. (2021a) for some evaluations of the reviewer assignment procedures in computer science.
After initial reviews are submitted by the reviewers, authors get a chance to rebut any incorrect statements in the reviews (Gao et al., 2019; Liu et al., 2023). The reviewers get to read the rebuttal, optionally update their reviews, and engage in a discussion with each other on a typed forum (Stelmakh et al., 2020; Rastogi et al., 2024). A meta-reviewer (equivalent of an associate editor in journals) is also involved in the discussion who eventually makes a recommendation to the program chair on which papers should be accepted. Typically, about 20%–25% of the submitted papers are accepted to the conference.
3.4 Influence on ALMA
HST, JWST, ESO, and computer science have shown the viability of machine learning and optimization techniques that inspired ALMA’s efforts. ALMA’s implementation differs from those used by other observatories in a few key ways. First, the reviewer’s expertise was established based on their ALMA proposal history rather than ADS data. Each reviewer is guaranteed to be on at least one ALMA proposal by design of the system, and a reviewer’s proposals can be identified unambiguously. Second, proposal categories and keywords are combined with reviewers’ expertise keywords to refine assignment options, reducing the number of spurious matches that machine learning alone might generate. Lastly, we adopted a computer science optimization algorithm for assigning proposals, with a few adjustments to fit ALMA’s distributed peer review model. Section 4 outlines the deployment of this new assignment process in detail.
4 Deployment
This section describes the deployment of the proposal assignment process adopted by ALMA and implemented for 2023 (Cycle 10). Figure 1 provides an overview of the process from the proposal submission by the PI to the ranking of proposals by a reviewer. The submitted proposal is passed to unsupervised machine-learning algorithms to infer the topics of the proposal (see Section 4.1). The PI designates a reviewer that will participate in the review process, where the scientific expertise of the reviewer is established from their proposal history, also based on machine learning techniques (see Section 4.2). The similarities between the proposal topics and the reviewer’s expertise is computed, and after flagging conflicts and disallowed proposal assignments, the proposals are assigned to the reviewers (see Section 4.3). The following subsections describe the deployment in more detail.
Figure 1: Flowchart of the proposal assignment process used by ALMA in Cycle 10, starting from the proposal submission by the PI and ending with the reviewer providing their rankings and self-assessment of their expertise on individual assignments.
4.1 Topic Modeling
The first step in the proposal assignment process is to determine the topics of each proposal using natural language processing. The PDF files for submitted ALMA proposals from Cycle 1 onwards were converted to ASCII text using the python library pdftotext. [^3] https://github.com/jalan/pdftotext The technical justification portion of the proposal that describes and justifies the setup of the observations is removed given it is largely technical in nature and much of the text will appear similar between proposals. The ASCII text for the scientific justification was then combined with the proposal title and proposal abstract stored in the ALMA databases to define the proposal text. Pre-processing steps were applied to the resulting files using spaCy (Honnibal et al., 2020) and custom functions to define and remove stop words, standardize commonly used acronyms, convert British spelling of words to American spelling, and lemmitization. This process created a “bag of words” for each proposal.
Many methods are available to classify documents from the bag of words. We mainly explored the use of TF-IDF, Latent Semantic Indexing, and Latent Dirichlet Allocation (LDA). All three methods gave satisfactory and similar results based on how well they can find similar documents which are known apriori to share the same scientific category and keywords. We ultimately settled on the LDA method (Blei et al., 2003) since we found for our application that it was able to match proposals to within the same scientific category at a slightly higher rate than TF-IDF and LSA.
LDA is an unsupervised machine learning algorithm that assumes a collection of documents can be represented by a mixture of topics, and each topic can be represented by a mixture of words. For a specified number of $N$ topics, the LDA algorithm processes a collection of proposals to define the $N$ topics, and computes a vector for each proposal that contains the relative weight of each topic for that proposal. Thus each proposal $i$ is represented by an $N$-length vector $\vec{p_{i}}$ that indicates the relative strength of the $N$ topics. The topics were identified using the 17,051 proposals submitted to ALMA starting in Cycle 1 and up to and including Cycle 10. We used the LDA algorithm implemented within the python package gensim. The number of topics is a free parameter that was determined upon iteration by inspecting the words of the topics to judge if the topics were sufficiently distinct based on the knowledge of the type of proposals submitted to ALMA. In addition, we identified the most similar proposals to a given proposal to investigate how well proposals from the same scientific category and keywords can be matched. While subjective, we settled upon $N=50$ topics.
To gauge the efficacy of our use of the LDA algorithm, we computed the similarity between any two proposals $i$ and $j$ as
$S(\vec{p_{i}},\vec{p_{j}})=\frac{\vec{p_{i}}\cdot\vec{p_{j}}}{||\vec{p_{i}}||_{2}\,\,||\vec{p_{j}}||_{2}}.$ (1)
Here, $S(\vec{p_{i}},\vec{p_{j}})$ ranges between 0 and 1, where a value of 0 indicates that the topics of the two proposals are dissimilar and a value of 1 indicates the topics are the same. For each proposal, the 10 proposals that have the highest similarity were identified, and we computed the fraction of those proposals that were submitted to the same scientific category as selected by the PI. While this is by no means a perfect measure of the accuracy of the algorithm since there are ambiguities in the definitions of the scientific categories, it does provide a first order indication of how well the algorithm performs since proposals in the same scientific category are expected to have similar content.
Figure 2 shows the results for the Cycle 10 proposals. As an example, for proposals submitted to Category 1, 80% of the matched proposals were also in Category 1, 17% in Category 2, 2% in Category 3, 0% in Category 4, and 1% in Category 5. The high overlap between Categories 1 and 2 is expected given both contain galaxy proposals and there is no precise division by redshift. More relevant is that only 3% of the proposals were matched in Categories 3, 4, and 5 combined. Even the 1% matching in Category 5 is not necessarily spurious, since this category contains transient proposals which overlaps with Gamma Ray Bursts in Category 1. Category 5 is the one category that has significant overlap with all other categories. This reflects that large range of topics in this category (e.g., the Sun, transients, pulsars, evolved stars, blackholes), and it also contains the fewest overall proposals, so there may not be 10 proposals that are closely matched. The main conclusion is that while most of closest matched proposals are in the same category as the submitted proposal, there are potential false matches.
Figure 2: Comparison of scientific categories for the closest matching proposals identified using the Latent Dirichlet Allocation model, alongside the category of the submitted proposal. The x-axis represents the categories of the submitted proposals, while the y-axis displays the categories of the ten most similar proposals. Each column is normalized to unity. The results are shown for proposals submitted to ALMA Cycle 10. For each submitted category, the closest matching proposals identified by the machine learning model are found within the same category.
4.2 Reviewer expertise
The PI of a proposal must specify one person from the proposal team that will serve as a reviewer in the review process. Most PIs serve as the reviewer on behalf of their proposals, but in about 12% of cases, a co-Investigator (coI) is selected (Donovan Meyer et al., 2022). The reviewer must be a member of the proposal team and registered within the ALMA system.
In previous applications of machine learning to proposal assignments at astronomical observatories (Kerzendorf, 2019; Kerzendorf et al., 2020; Strolger et al., 2023), the reviewer expertise has been inferred based on published abstracts or papers that are available in ADS. This approach has the advantage in that it contains an extensive record of a reviewer’s publication history and therefore expertise. While the publication record of most authors can be inferred accurately from their last name and first initial (Milojević, 2013), erroneous matches will inevitably occur. With over 1000 reviewers participating in the review process, it is not practical to perform a manual inspection of the search results to identifying any problematic identifications. While an Open Researcher and Contributor ID (ORCID) can be used to identify authors uniquely, not all papers are tagged with an ORCID, [^4] https://orcid.org and ALMA does not currently collect ORCIDs from its users. Also, ALMA allows students who are the PI of proposals to serve as reviewers by specifying a mentor, and students may not yet have a publication record.
Instead of using the publication record, ALMA used the proposal history of the reviewers, including the current proposal cycle, to establish their scientific expertise. The main advantage of this approach is that reviewers can be uniquely identified by their ALMA user ID, and by design of the review process, each reviewer is guaranteed to be an investigator for at least one proposal. In addition, each proposal is tagged by the proposal PI with a scientific category and keywords that provides additional information on the topic of the proposal. As discussed below, these keywords can help identify which proposals should be used to determine the reviewer’s expertise and to prevent proposal assignments in disparate scientific categories (see Section 4.3). A downside to using the proposal text is that the proposal history will not capture the full expertise of a reviewer, especially for senior astronomers who are new users to ALMA and may have a substantial publication history.
Each reviewer is associated with one or more proposals as either PI or coI. The topical vector for each proposal associated with a reviewer is determined as described in Section 4.1. These topical vectors are averaged to determine the reviewer’s overall expertise ($\vec{r}$), which is also an $N$ length vector. Similar to Equation 1, the cosine similarity between the $i^{th}$ proposal and the $k^{th}$ reviewer’s expertise is used to assess where the reviewer is suitable for evaluating that proposal:
$S(\vec{p_{i}},\vec{r_{k}})=\frac{\vec{p_{i}}\cdot\vec{r_{k}}}{||\vec{p_{i}}||_{2}\,\,||\vec{r_{k}}||_{2}}.$ (2)
The similarity score $S(\vec{p_{i}},\vec{r_{k}})$ ranges from 0 to 1, where 0 indicates no alignment between the reviewer’s expertise and the proposal topic, and 1 indicates a perfect match. While the absolute value of the similarity score may not be crucial, the relative values should provide a meaningful comparison of how well each proposal aligns with the reviewer’s expertise.
This general approach works well for most reviewers, but the implementation also accounted for those with expertise or proposal submissions across multiple categories. Special considerations were also made for categories and keywords with limited available reviewers. The Appendix provides further details on how reviewer expertise was calculated to address these specific situations.
4.3 Proposal assignments
Given the formalism to compute the similarities between the proposal topics and reviewer’s expertise as defined in Section 4.2, we proceed to assign proposals to reviewers. The goal is for each proposal to be reviewed 10 times and for each reviewer to be assigned 10 proposals. The main input is a matrix that contains similarities between the reviewer expertise and all submitted proposals for a given cycle. The rows in the matrix correspond to each submitted proposal, and the columns in the matrix correspond to each submitted proposal for the cycle. If an individual is reviewing on behalf of more than one proposal, they will appear as separate rows in the matrix.
Before assigning proposals, similarities are set to a negative number if a proposal assignment is not permitted. There are two main reasons to restrict proposal assignments in this manner. First, potential assignments are flagged if the reviewer has a known conflict of interest on the proposal. Conflicts of interest are defined if the reviewer is a PI or coI on the potential proposal assignment proposal, the PI is deemed a close collaborator based on the proposal history (see Donovan Meyer et al., 2022), or the reviewer identified a PI or coI as a close-collaborator. In these cases, the similarities are set to a value of $-100$, which effectively prohibits the assignment of the conflicted proposal. The second case is if the potential assigned proposal is from a distinct scientific category compared to the reviewer’s expertise (see Section 4.2). While such proposals will typically have low similarities, this additional measure reduces the risk of a false positive match. In practice, the new similarity is set as $-1.01+S(\vec{p_{i}},\vec{r_{k}})$. While a negative number does not preclude a proposal assignment, it makes it less likely and can be checked after the assignments are made.
Once the similarity matrix is established, proposals are assigned to reviewers. The goal is to define a metric that measures the success of these assignments and optimize it during the assignment process. One common approach is to maximize the total similarity scores between proposals and reviewers, which increases overall expertise in the review process. However, this method may lead to unfairness if certain types of proposals are disproportionately assigned to less-qualified reviewers. ALMA instead adopted the approach used in PeerReview4All, which optimizes a “leximin fairness” metric (Stelmakh et al., 2021a, code available at https://github.com/niharshah/peerreview4all). This method optimizes the similarity for the most “disadvantaged” proposals; i.e., proposals that have the least overall expertise among the reviewers. Once the most disadvantaged proposal’s assignment is optimized, the algorithm moves to the next most disadvantaged, and so on. Proposals in fields with many qualified reviewers are assigned last, as they have more reviewer options available. The metric used by PeerReview4All is solved using the gurobi python optimizer (Gurobi Optimization, LLC, 2023).
The PeerReview4All code was adapted to handle special cases. Student reviewers specify a mentor, who may also serve as a reviewer for other proposals. To ensure fairness, we required that the proposal sets for students and mentors be disjoint. After each round of assignments, the similarity matrix was updated to mark as conflicted any proposals previously assigned to a reviewer or their mentor. Additionally, conflicts were updated for reviewers assigned multiple sets, ensuring they were not assigned the same proposal twice.
To prevent “niche” proposals from repeatedly being reviewed by the same group, further restrictions were imposed on the number of proposals a reviewer could be assigned within the same topic. These restrictions specifically applied to proposals related to Very Long Baseline Interferometry, Target of Opportunity, and solar studies.
5 Evaluation
Given the framework for computing the similarity between proposals and reviewer’s expertise described in Section 4, we now evaluate if the new process improved the proposal assignments to the reviewers compared to previous cycles. Recall that the ALMA reviewing cycle where the new assignment approach was implemented is Cycle 10, and the two previous cycles we use for backtesting are Cycle 8 and Cycle 9.
5.1 Backtesting similarity computation on Cycle 8 and Cycle 9 data
We backtest the effectiveness of the similarity scores (calculated as described in Section 4.2) by using data from previous review cycles. Specifically, we examine whether the computed similarities between a reviewer’s expertise and the proposal topics accurately reflect the reviewer’s expertise. During the distributed peer review process used by ALMA in Cycles 8 and 9, reviewers were asked to self-assess their expertise on each assigned proposal by indicating if the assigned proposal is in their field of expertise (i.e., the review considers themselves an “expert”), they are somewhat knowledgeable, or they have little to no knowledge of the topic. If our similarity scores are a reliable measure of expertise, we would expect that reviewers who self-identify as experts in a proposal topic will, on average, have higher similarity scores compared to those who report having little or no knowledge of the topic.
Figure 3 shows the similarities of the proposal assignments in Cycle 8 and 9 based on the reviewer’s self-declared expertise. In these cycles, assignments were made solely using categories and keywords, without considering similarity metrics. When reviewers identified a proposal as being within their area of expertise, the distribution of similarities has a median of approximately 0.35 for both cycles. In contrast, when reviewers indicated only partial familiarity with a proposal, the median similarity dropped to around 0.12, and for those with little to no expertise, the median similarity was 0.04.
As expected, the highest median similarity occurred when reviewers declared expertise, and the lowest when they indicated little or no knowledge, with comparable trends across Cycles 8 and 9. However, similarity alone is not a perfect indicator of expertise. Some assignments had near-zero similarity scores, yet reviewers reported they were experts, while other proposals had similarity scores close to 1, but reviewers expressed limited or no knowledge of the subject. This may point to both the inherent imprecision in survey data and the limitations of assessing expertise based solely on proposal history. Despite these limitations, the correlation between self-declared expertise and similarity suggests that these metrics can help improve the alignment of proposals with a reviewer’s scientific background.
Figure 3: Cumulative distributions of the similarities between the reviewer’s expertise and the topic of the assigned proposal in Cycle 8 (left) and Cycle 9 (right), when categories and keywords were used to assigned proposals. The similarities were computed retroactively to backtest the machine learning algorithm. The results are shown for proposal assignments where reviewers indicated they are experts (solid curve), had some knowledge of the proposal (dashed), and had little/no knowledge of the proposal (dotted). Similarities tend to be higher when reviewers indicate expertise in the proposal, suggesting that these similarities have predictive value for assessing the suitability of reviewer assignments.
5.2 Comparing Cycle 10 and Cycle 9 assignments, given computed similarities
In this section, we analyze the role of the optimization procedure for computing the assignments (Section 4.3) given the computed similarities. In Cycle 9, proposals were assigned to reviewers based on the proposal category and keywords (Donovan Meyer et al., 2022); i.e., the “old” algorithm. In Cycle 10, we used the similarities and assignment algorithm described in Section 4; i.e., the “new” algorithm. In order to make a fair comparison, we recompute similarities for Cycle 9 using the software developed for Cycle 10. In the analysis in this subsection, we use these similarities. In both cycles, a small percentage of the initial assignments were reassigned after reviewers declared conflicts of interest. [^5] The percentage of assignments where reviewers declared a conflict of interest increased in Cycle 10 to 2.4% from 1.4% in Cycle 9, as more reviewers declared the assignments conflicted scientifically with their own proposal when the machine learning algorithm was adopted.
Figure 4 shows a histogram of the similarities in the actual proposals assignments for Cycle 9 (blue histogram), and Cycle 10 (orange histogram) which used the optimization method of Section 4.3. We find a substantial increase in the similarity between the reviewer expertise and the assigned proposals compared to the previous algorithm, with the median similarity increase to 0.71 in Cycle 10 from 0.20 in Cycle 9.
Figure 4: Histogram of the similarities between the reviewer expertise and the assigned proposal for Cycle 9 under the old algorithm (blue, hatched histogram), and in Cycle 10 for the new algorithm (orange, dotted histogram). With the new assignment process adopted in Cycle 10, the median similarity increased to 0.71 from 0.20 in Cycle 9, indicating a higher level of expertise in the proposal assignments was achieved with the new assignment process (see Section 4.3).
5.3 Cycle 10 reviewer survey
The increased similarity of the proposal assignments was also reflected in the reviewer survey, where reviewers provided a self-assessment of their expertise on each proposal assignment. Figure 5 shows a histogram of the results from Cycles 8, 9 and 10. With the new assignment process adopted in Cycle 10, the percentage of reviewers indicating they were an expert in the assigned proposal increase to 65% in Cycle 10 from 45% in Cycle 9. Similarly, the percentage of assignments where the reviewers indicated they had little or no knowledge on the proposal declined to 5% in Cycle 10 from 10% in Cycle 9.
Figure 5: Histogram of the reviewer’s self-expertise on their proposal assignments in Cycles 8–10. With the new assignment algorithm implemented in Cycle 10, the percentage of reviewers reporting expertise in their assigned proposals increased by 20 percentage points compared to the previous cycle, while the percentage of those declaring little or no knowledge was reduced by half.
5.4 Topological accuracy of the proposal rankings
In this section, we evaluate a key tradeoff in expertise-based reviewer assignments: balancing the expertise of assigned reviewers with the topology of the assignments. Reviewer expertise is crucial for evaluating research proposals, prompting strategies that prioritize maximizing relevant expertise. However, strictly focusing on expertise can lead to the isolation of certain proposal groups, especially those within the same subtopic. For example, imagine a scenario where half of the reviewers are exclusively assigned to review one group of proposals, while the other half reviews the rest. This division hampers direct comparison between the two groups, limiting the ability to rank proposals effectively across groups.
To quantify how assignment topology affects ranking accuracy, we use the methodology from Shah et al. (2016) to assess how the structure of reviewer assignments impacts the final outcomes, assuming all reviewers have equal expertise. Consider a graph where each proposal is a vertex. An edge is placed between two vertices if there is at least one reviewer who reviews both proposals (that is, if the two proposals are directly compared). The weight of an edge is the number of times that pair of proposals is assigned to the same proposal set. The Laplacian for this graph determines the topology of the proposal assignments. Following Shah et al. (2016, see also , ), the trace of the (pseudo) inverse of the Laplacian is proportional to the ranking accuracy.
To estimate the relative accuracy of the actual assignments in Cycle 10, we set the error in the actual assignments to unity. We then ran simulations where the reviewers were assigned proposals at random across any scientific category, subject to the constraint that each reviewer is assigned 10 proposals and each proposal is assigned to 10 reviewers. The random assignment strategy is known to be optimal (Shah et al., 2016) in terms of topology when all reviewers have identical expertise. We find that the relative uncertainty is 0.86 for the random assignment. Thus, optimizing assignments for expertise increases the uncertainty in final outcomes by only 14% based on topology of the proposal assignments.
6 Discussion and conclusions
The implementation of the new assignment process in ALMA Cycle 10, leveraging machine learning (Blei et al., 2003) and optimization (Stelmakh et al., 2021a) algorithms, has successfully met the high-level objectives set for the system. The similarity measures between the proposal topics and reviewer’s expertise computed from their proposal history are correlated with the reviewer’s self-reported expertise to demonstrate the similarity measures have predictive value. With the new optimization algorithm, the average similarity is significantly higher than in previous cycles with the old algorithm. The success of the algorithm is confirmed in the Cycle 10 reviewer survey, which showed that reviewers reported a higher level of expertise for the assignment than the previous two cycles. Finally, the initial assignments were run without any manual adjustments, resulting in significant time savings in handling the proposals and indicating the robustness of the process.
Our rigorous evaluation demonstrates the potential for widespread adoption of AI-driven reviewer assignments in both proposal and paper reviewing, offering substantial improvements. This methodology is especially useful for scientific disciplines that continue to struggle with manual assignment processes, where scalability and growing complexity demand more sophisticated, automated approaches.
6.1 Limitations
We now discuss some limitations of our study. Firstly, a portion of our findings is based on reviewers’ self-reported expertise regarding the proposals they assessed. Such self-reports are susceptible to various biases, suggesting that future research should incorporate additional evaluative metrics. Secondly, our analysis largely compares the current cycle to previous ones. There is a possibility that other significant changes, such as shifts in the types of proposals or the reviewer demographic, could have influenced the results. To counteract potential confounding factors, we conducted an analysis using the current algorithms on data from previous cycles, which affirmed the consistency of our results. Additionally, the firsthand experience of the first two authors, who have overseen this review process for several cycles, indicates no major changes during this period.
Our experience found that while the machine learning algorithms were quite successful in identifying similar proposals, a small fraction of assignments were matching proposals triggered on some individual words that are relatively rare, but can appear across different scientific categories (e.g., polarization, maser, corona). Some of these mismatches could likely be reduced with improved machine learning algorithms (e.g., using transformer-based language models and/or fine tuning to ALMA-relevant corpora), but the possibility of mismatches will always remain. In the current deployment, by using the keywords of the proposals and reviewer expertise, inappropriate assignments were flagged in advance and increased the robustness of the overall process.
6.2 Open problems
Finally, we outline several open problems in this exciting research domain. First, in assigning reviewers to the proposals, there is a tension between expertise and topology of assigned reviewers, as discussed in Section 5.4. While previous works have explored these factors independently, a promising direction is to study them jointly and establish the optimal tradeoff.
Second, unlike many other review systems that rely on ratings, ALMA requires each reviewer to rank the proposals they evaluate. Determining the most effective evaluation method is a key area of interest, closely tied to the reviewer assignment process. Rankings present challenges for cross-reviewer comparisons due to their independence, whereas ratings are subject to calibration issues such as varying strictness or leniency across reviewers (Ge et al., 2013; Roos et al., 2012; Wang & Shah, 2019). Additionally, research is ongoing into methods for eliciting both rankings and ratings and effectively integrating them (Pearce & Erosheva, 2022; Liu et al., 2022).
Third, systems like ALMA accumulate multiple cycles of reviews, presenting an opportunity to develop methods that leverage these data to continually improve reviewer assignments.
Fourth, competitive reviewing scenarios may incentivize strategic or dishonest behavior. One example is the formation of “collusion ring”, where reviewers support a paper in exchange for reciprocal favors (Littman, 2021; Vijaykumar, 2020). Colluding reviewers may manipulate the assignment system to be assigned their target paper. Some venues mitigate this risk by introducing randomization into the assignment process (Jecmen et al., 2020), limiting the likelihood of such matches. While colluders may still achieve high similarity scores, randomization reduces the probability of assignment. Another issue arises when reviewers deliberately downrate competing papers to increase the chances of their own being accepted (Balietti et al., 2016). Research in this area focuses on developing optimization methods that prevent reviewers from influencing the outcomes of their own papers (Alon et al., 2011; Dhull et al., 2022) and detecting large-scale manipulative behavior (Stelmakh et al., 2021b). Improving detection and mitigation of such behavior remains an open challenge.
As ALMA continues to refine its assignment process in future cycles, the successful deployment of AI-based methods underscores the potential for broader adoption of these technologies across the scientific community, offering a powerful tool to address both current and future challenges in peer review.
Acknowledgements
We thank Victoria Catán and Nicolas Buzeta for their help in developing the initial topic modeling algorithms, and Ignacio Toledo for his assistance in using the Dataiku platform. The work of NBS was supported by NSF CAREER 1942124.